OpenStudy (anonymous):

What is the y-intercept of the line whose equation is y =1/2 x - 3? -3 1/2 3

OpenStudy (anonymous):

y=ax+b b=y-intercept

OpenStudy (anonymous):

thanls =P

OpenStudy (anonymous):

-3

OpenStudy (anonymous):

why is it -???

OpenStudy (anonymous):

From the formula y = mx + b where b is the y-intercept, compare it with your equation y = mx + b y = 0.5x - 3 b = -3 -> y intercept

OpenStudy (anonymous):

b=-3